Derive Clausius-Clapeyron equation for the
first-order phase transition.
1
Expert's answer
2019-05-07T09:56:19-0400

Two phases are in equilibrium when

"\mu_1(T,P)=\mu_2(T,P)"

So

"d\mu_1(T,P)=d\mu_2(T,P)"

"\left(\frac{\partial \mu_1}{\partial T}\right)_P dT+\left(\frac{\partial \mu_1}{\partial P}\right)_T dP=\left(\frac{\partial \mu_2}{\partial T}\right)_P dT+\left(\frac{\partial \mu_2}{\partial P}\right)_T dP"

Using Maxwell's relations

"\left(\frac{\partial \mu}{\partial T}\right)_P=-S"

"\left(\frac{\partial \mu}{\partial P}\right)_T =V"

we obtain

"-S_1dT+V_1dP=-S_2 dT+V_2 dP"

Finally

"\frac{dP}{dT}=\frac{S_2-S_1}{V_2-V_1}"

Since

"T\Delta S=Q"

we also have

"\frac{dP}{dT}=\frac{Q}{T(V_2-V_1)}"